1. If your car has a 20. gallon gas tank full of octane (C8H18) how much CO2 is produced in grams if you have complete combustion of all 20. gallons? (Density of octane = 0.702 g/mL)
1
Expert's answer
2020-04-13T02:25:33-0400

1 gallon=3.8 litres

20 gallons=20"\times"3.8 litres =76 litres

given that density of octane is 0.702 g/ml

hence mass of 76( or 76000 ml) litres of octane = 0.702"\times"76000=53352 grams

molar mass of octane = 114 g

hence number of moles of octane ="\frac{53352}{114}" =468moles


"C_8H_{18}(g)+12.5O_2(g)\u21928CO_2(g)+9H_2O(g)"

so from the reaction we can see that one mole of octane produces 8 mole of carbondioxide

hence 468 moles will produce (468"\times"8) moles of carbondioxide

3744 moles of CO2 is produced

mass of 3744 moles of carbondioxide is "3744\times44=164736 grams"